OpenStudy (anonymous):

The function T(n) is defined below. Which of the following are equal to T(8)? Check all that apply. T(n) = 4n - 5 A. 27 B. 37 C. T(5) + T(3) D. T(7) + 4

OpenStudy (hba):

Just plug in 8 in place of n :)

OpenStudy (anonymous):

T(n) = 4n - 5 for T(8) just replace n = 8 in the equation

OpenStudy (anonymous):

t(8) = 4* 8 - 5 = 27

OpenStudy (anonymous):

.... i hate my life ........